Smokies BBQ is considering raising their prices. Currently, their cost per BBQ sandwich is $2.44 and they sell the sandwiches based on achieving a margin of 40%. They sell 427 sandwiches each day. If they decide to use a 61% margin instead, how much will their total contribution increase per day if they sell the same number of sandwiches at the new price?
Answer #1

Solution:

Initial Scenario:

Cost per BBQ Sandwich = $2.44

Margin = 40% of CP = 0.4*2.44 = $ 0.976

No. of sandwiches sold each day = 427

Therefore, total contribution = 427 * 0.976 = $ 416.752

Proposed Scenario:

New Margin = 61% of CP = 0.61*2.44 = $1.4884

New total contribution for same no. of sandwiches i.e 427 = 427 * 1.4884 = $ 635.5468

Increase in total contribution = New contribution - Initial Contribution = 635.5468 - 416.752 = $ 218.7948

Ans: Total contribution will increase by $ 218.7948